### v2

# Towards Robust Infrared Small Target Detection: A Feature-Enhanced and Sensitivity-Tunable Framework

Abstract
Recently, single-frame infrared small target (SIRST) detection technology has attracted wide-spread attention. However, due to the intrinsic feature scarcity in infrared small targets, precise segmentation of small targets from complex backgrounds remains a significant challenge. Different from most existing deep learning-based methods that focus on improving network architectures, we propose a feature-enhanced and sensitivity-tunable (FEST) framework, which is compatible with existing SIRST detection networks and further enhances their detection performance. The FEST framework improves the model’s robustness from two aspects: feature enhancement and target confidence regulation. For feature enhancement, on the one hand, we adopt a multi-scale fusion strategy, which can effectively improve the model’s perception and adaptability to multi-scale features of multi-size targets. On the other hand, we construct an edge enhancement difficulty mining (EEDM) loss based on the analysis of the task characteristics, which helps guide the network to continuously focus on challenging target regions and edge features during training. For target confidence regulation, we design an adjustable sensitivity (AS) strategy for network post-processing. This strategy not only enhances the adaptability of the network in complex scenarios, but also significantly improves the detection rate of infrared small targets while maintaining segmentation accuracy. Extensive experimental results show that our FEST framework can significantly enhance the performance of existing SIRST detection networks. Notably, the multi-scale direction-aware network (MSDA-Net) equipped with the FEST framework won the first prize in the “PRCV 2024 wide-area infrared small target detection competition”.
keywords:
Infrared small target detection , Multi-scale fusion strategy , Adjustable sensitivity strategy , Edge enhancement difficulty mining loss
journal: *****************
[…]
1 Introduction
Infrared imaging technology, with its excellent low-light imaging capabilities, demonstrates unique advantages in complex environment perception, driving the widespread application of infrared detection technologies [1, 2, 3, 4]. Among them, SIRST detection, as a fundamental technology in infrared detection technology, plays an important role in many fields, including target recognition and tracking [5, 6], early warning systems [7, 8], traffic surveillance [9, 10], and maritime search and rescue [11, 12].
The research on SIRST detection can be divided into non-deep learning-based methods and deep learning-based methods. Non-deep learning-based SIRST detection methods mainly use the characteristics of infrared small target images for detection [13, 14, 15, 16, 17, 18, 19, 20]. These methods often employ traditional image processing techniques, including filtering, sparse representation, low-rank decomposition, and background modeling, to enhance target features while reducing background interference. They are usually optimized for specific scenarios and can achieve good detection results under specific imaging conditions. However, owing to the complexity and variability of practical scenarios, these traditional methods often suffer from many false detections or even failure. With the rapid development of deep learning, a large number of researchers have focused on leveraging the powerful representation capabilities of deep neural networks to extract richer, more efficient, and more robust features, which have been widely applied across various research fields [21, 22, 23, 24]. In light of current technological trends and growing application demands, deep learning-based SIRST methods [25, 26, 27, 28, 29, 30, 31, 32, 33] have achieved remarkable progress. This type of method usually uses deep learning models such as convolutional neural network (CNN) [34] and visual transformers (ViT) [35] to automatically learn the spatial and texture features of the target from the data, thereby improving detection performance. However, most current methods mainly focus on optimizing the network structure while ignoring other aspects that may improve the detection performance. Therefore, it is of great significance to design an efficient and robust SIRST detection framework to enhance the overall performance of networks in complex environments.
Different from most existing deep learning-based methods that focus on improving network architectures, we propose a feature-enhanced and sensitivity-tunable (FEST) framework. The framework can be flexibly compatible with existing SIRST networks, thereby improving its robustness in complex scenarios. Specifically, in the aspect of the feature enhancement, a multi-scale fusion strategy is employed to improve the model’s perception and adaptability to multi-scale features of multi-size targets. By integrating information across multiple spatial scales, a comprehensive representation of multi-scale features of a single target is achieved, and adaptability to the scale variation in multiple targets is enhanced. In addition, to guide the network to pay more attention to challenging target regions and edge features during training, we propose an edge enhancement difficulty mining (EEDM) loss. This loss helps the model learn valuable features more effectively and greatly avoids the risk of small targets being submerged by the background. In the aspect of the target confidence regulation, to further enhance the adaptability of our framework to different scenarios, we propose an adjustable sensitivity (AS) strategy for post-processing. This strategy introduces the concepts of strong targets and weak targets. It can significantly improve the detection rate of infrared small targets while ensuring a small change in the false alarm rate and segmentation accuracy, thus having important practical application value. Extensive experimental results show that the proposed FEST framework can significantly enhance the performance of existing SIRST detection networks. The contributions of this manuscript can be summarized as follows:
  • 1.
    A feature-enhanced and sensitivity-tunable (FEST) framework is constructed, which can be flexibly compatible with the existing SIRST network and effectively enhances their detection performance.
  • 2.
    For feature enhancement, a multi-scale fusion strategy is adopted to improve the model’s perception and adaptability to multi-scale features of multi-size targets. In addition, an edge enhancement difficulty mining loss is constructed, which can guide the network to pay more attention to edge features and difficult target regions.
  • 3.
    For target confidence regulation, an adjustable sensitivity strategy is proposed, which deeply mines the generated probability mask information and innovatively defines the concepts of strong targets and weak targets. This strategy is conducive to significantly improving the detection rate while ensuring the false alarm rate and segmentation accuracy.
2 Related Work
2.1 Non-deep learning-based SIRST detection methods
Non-deep learning infrared small target detection methods can be roughly divided into three categories: background suppression-based methods, human visual system-based methods and image data structure-based methods. Among them, background suppression-based methods include spatial domain filtering (e.g., Max-Median filter [13] and two-dimensional minimum mean square error filter [14]) and transform domain filtering (e.g., dual-tree complex wavelet [15] and bidimensional empirical mode decomposition (BEMD) [16]). Spatial domain filtering methods are computationally efficient but suffer performance degradation in complex scenes due to the difficulty in predefining effective templates. In contrast, transform domain methods offer stronger background modeling capabilities but incur high computational costs, limiting their practical applicability. The human visual system-based methods simulate the human eye’s perception of local contrast to detect salient objects in images. Representative approaches include the local contrast measure (LCM) [17] and multi-scale patch-based contrast measure (MPCM) [18]. These methods can effectively suppress large-area bright background interference and improve target saliency. However, they still tend to produce many false detections when strong noise or interference is present in the image. Image data structure-based methods formulate the detection task as a low-rank and sparse matrix separation problem, leveraging the non-local self-similarity of the background and the sparsity of the target to achieve foreground-background separation. Representative approaches include the infrared patch-image model (IPI) [19] and the non-negative infrared patch-image model based on partial sum minimization of singular values (NIPPS) [20]. These methods generally offer effective separation performance but remain sensitive to strong edges in the background and are computationally expensive.
Figure 1: Overall structure of the FEST framework. SIRST Network denotes the selected existing SIRST networks. Size $O_{1}$, …, Size $O_{n-1}$, and Size $O_{n}$ denote the actual sizes of the original training images. Size 1, Size 2, and Size 3 denote the input resolutions corresponding to each model. Size $O_{t1}$, …, Size $O_{t(m-1)}$, and Size $O_{tm}$ denote the actual sizes of the original test images.
2.2 Deep learning-based SIRST detection methods
Different from non-deep learning-based methods, deep learning-based methods learn the features of infrared small targets in a data-driven manner [36, 37]. Specifically, Dai et al. propose the asymmetric context module (ACM) [25] and the attention local contrast network (ALCNet) [26], which for the first time introduced the local contrast idea in traditional methods into the deep learning-based method. To further explore the relationship between local areas, Yu et al. successively proposed the multi-scale local contrastive learning network (MLCL-Net) [27] and the improved local contrastive learning network (ALCL-Net) [28], which enhanced the network’s perception ability of infrared small targets. Subsequently, Li et al. proposed a densely nested attention network (DNANet) [29] based on the UNet++ architecture , aiming to alleviate the problem of target information loss caused by downsampling operations in infrared small target detection. At the same time, Wu et al. proposed a U-Net nested network (UIU-Net) [30], which achieves richer multi-scale feature expression by nesting a lightweight U-Net network in the backbone U-Net. In recent years, Zhao et al. further emphasize the gradient information and directional characteristics of infrared small target images and propose GGL-Net [31] and MSDA-Net [32] respectively. GGL-Net introduces amplitude images at multiple stages of network training, prompting the network to pay more attention to discriminative features. MSDA-Net further promotes the refinement of infrared small target segmentation by emphasizing the attention to multi-directional high-frequency features. In addition, Zhang et al. proposed an infrared shape network (ISNet) [33] to improve the accuracy of target recognition by emphasizing the extraction of shape features in infrared images. However, most current methods focus on optimizing the network structure while ignoring other aspects that may improve the detection performance. Therefore, we aim to build an efficient and robust SIRST detection framework to further improve the performance of existing SIRST detection networks.
3 Method
3.1 Overall Framework
From fig. 1, the proposed scheme consists of two parts: the model training phase and the model inference phase. In the model training phase, first, we perform multi-scale transformations on the original training images, adjusting them to multiple resolutions through scaling operations. Then, the images at different scales are individually fed into the corresponding basic models for training. The basic models can be any existing SIRST network. Finally, to further improve the detection accuracy of the network, we use the proposed EEDM Loss as the loss function to optimize the network. In the model inference phase, first, we perform multi-scale transformations on the original training images and input the transformed images into the corresponding multi-scale models for inference. Next, the feature maps generated by multiple scale models are uniformly resampled to the original test image size to ensure feature alignment. Then, the multi-scale fusion strategy is used to integrate the prediction results of multiple scale models. This strategy can improve the model’s perception of different target scales and enhance its robustness in complex environments. Finally, the proposed AS strategy is used to post-process the fused results. This strategy can significantly improve the detection rate of infrared small targets while ensuring the false alarm rate and segmentation accuracy.
3.2 Multi-scale fusion strategy
In the SIRST detection task, the target scale is affected by factors such as imaging scene, shooting distance and sensor equipment, and can usually be divided into point targets, spot targets and extended targets [4, 32]. In complex scenes, a single-scale network model struggles to handle small targets of various sizes simultaneously, often resulting in missed or false detections. Additionally, considering that this study aims to build a detection framework compatible with the existing SIRST network, we seek to enhance detection performance without altering the network structure. Therefore, we adopt a multi-scale fusion strategy to enhance the network’s adaptability to different scales targets by integrating feature information from multiple scale models, effectively alleviating the shortcomings of a single-scale model.
From fig. 1, the multi-scale fusion strategy is reflected in three parts. First, in the model training phase, without modifying the structure of the SIRST network, images of different sizes are used for training to obtain models suitable for different resolutions. Secondly, in the inference phase, the original input image is transformed into multiple sizes and fed into the corresponding models for inference. Finally, to fully leverage the detection advantages of models at different sizes, the prediction results are resampled back to the original image size and integrated to produce the detection result.
Figure 2: Adjustable sensitivity strategy. Each connected component in the figure denotes an infrared small target. Green denotes correctly included weak targets, while red denotes incorrectly included weak targets.
The multi-scale model fusion strategy is conducive to fully mining the semantic features and detail features of multi-size targets at different scales, making the model more accurate and reliable in identifying and locating targets. In addition, practical detection scenarios are often more complex and changeable, so fusing models of different scales can help improve the mode’s robustness in complex backgrounds. In practical applications, the fusion scales and the number of models can be reasonably chosen according to specific task requirements to achieve an optimal balance between detection performance and inference efficiency.
3.3 Edge enhancement difficulty mining loss
In the task of SIRST detection, the small target area and lack of intrinsic features make it difficult to accurately locate in the image. Existing studies have shown that paying attention to edge information in this task can bring significant performance improvements [31, 33]. Therefore, on the basis of the characteristics of infrared small targets, we design an EEDM loss. This loss can effectively guides the network to pay more attention to edge features and challenging target regions during training, thereby enhancing the model’s boundary awareness and improving its ability to learn from challenging samples.
The EEDM loss consists of two components: edge pixel enhancement and difficult pixel mining. Edge pixel enhancement is performed preferentially. Specifically, the edge contour of the infrared small target is first extracted from the truth label obtain the edge map $e$, which is then weighted using corresponding edge weight coefficient $w$. Secondly, the binary cross-entropy loss [38] is employed to compute the loss value for each pixel, thereby constructing the matrix ${L_{BCE}}$. Finally, the edge weighting matrix is used to apply a weighted calculation to the loss matrix ${L_{BCE}}$, resulting in the weighted loss value ${L^{*}}$. The formulas are as follows:
${L_{BCE}}={y_{i}}\log({\hat{y}_{i}})+(1-{y_{i}})\log(1-{\hat{y}_{i}})$ (1)
[…]
${L^{*}}=e^{\prime}\times{L_{BCE}}$ (5)
where ${y_{i}}$ denotes the true label of the $i_{th}$ pixel, ${\hat{y}_{i}}$ denotes the probability of the $i_{th}$ pixel being predicted as the target, and $e^{\prime}$ denotes the edge weighting matrix.
The difficult pixel mining is performed subsequently. Specifically, first, the loss values corresponding to each pixel in ${L^{*}}$ are sorted. Secondly, pixels with loss values greater than or equal to the difficult pixel mining ratio are selected as the set of difficult pixels ${L_{{\rm{hard}}}}$. Finally, the average loss of the difficult pixels is computed to obtain the final loss value. The formulas are as follows:
${L_{sorted}}=sort({L^{*}})$ (6)
${L_{{\rm{hard}}}}=\left\{{{L_{i}}\mid{L_{i}}\geq{L_{{\rm{sorted}}}}\left[{\left\lfloor{p\times N}\right\rfloor}\right]}\right\}$ (7)
${L_{{\rm{EEDM}}}}=\frac{1}{{|{L_{{\rm{hard}}}}|}}\sum\limits_{i\in{L_{{\rm{hard}}}}}{{L_{hard}}}[i]$ (8)
where $p$ denotes the difficult pixel mining ratiol and ${L_{EEDM}}$ denotes the final output.
On the one hand, EEDM loss uses edge information as an additional constraint by assigning higher weights to the target boundary regions, thereby enhancing the model’s sensitivity to target boundaries. On the other hand, it dynamically discards a certain proportion of simple samples by difficult pixel mining, which enables the network to focus more on challenging regions and promotes the network to learn more discriminative features.
3.4 Adjustable sensitivity strategy
In the infrared small target detection task, the network outputs a probability mask, in which the value of each pixel denotes its confidence of belonging to the target. Existing studies typically use a single threshold (0.5) to separate targets from the background [4, 32]. However, we believe that the inherent idea of directly using a single threshold cannot fully mine and utilize the information in the probability mask. Therefore, we propose an AS strategy aimed at deeply mining the information in the probability mask. Based on this, we innovatively define the concepts of strong targets and weak targets. Strong targets refer to those that are easy for the network to detect and have high confidence. These targets typically display more distinct features in the image, enabling the network to quickly identify them through learning. For strong targets, our goal is to achieve complete and precise segmentation. In contrast, weak targets refer to those that are difficult for the network to detect and have lower confidence. These targets may be challenging for the network to recognize due to their shape or the complexity of the background. For weak targets, our goal is to identify them to meet the needs of practical scenarios.
Figure 3: Some samples from multiple datasets.
Through careful observation of the probability mask generated by deep learning methods in this task, we interestingly found that, influenced by data ambiguity and the network’s feature extraction mechanism, the network’s predicted probability mask for the target region exhibits a Gaussian-like distribution. Specifically, the probability is higher in the center region and gradually decreases towards the edge, and finally approaches zero. Based on this observation and to better detect strong and weak targets, we design two thresholds th1 and th2 ($th1>th2$) to adjust the model’s responsiveness to regions with different confidence levels. Specifically, when the network’s confidence in a region is higher than th1, the region is defined as a strong target region. In this case, the network demonstrates high detection capability and can provide relatively accurate segmentation results. Additionally, when the confidence is between th1 and th2, the region is defined as a weak target region. For these weak target regions, although the network’s detection accuracy is low, reasonable post-processing can still help identify partial features of these targets to meet the detection needs of weak target detection in practical scenarios.
The specific process of the AS strategy is shown in fig. 2, first, we separate the segmentation result ${M_{th2}}$ of th2 into multiple binary images $M_{th2}^{s},\;s=1,2,\ldots n$ with a single target region according to a single connected region. Secondly, these binary images $M_{th2}^{s},\;s=1,2,\ldots n$ are compared with the segmentation results ${M_{th1}}$ of th1 in turn to determine whether there are overlapping regions. For $M_{th2}^{s}$ with overlapping regions, we lose it. For the same target region, the fineness in ${M_{th1}}$ is better than that in ${M_{th2}}$. Thirdly, we extract the centroid of $M_{th2}^{s}$ where there is no overlapping region. Finally, the extracted centroid points are injected into ${M_{th1}}$ to generate the final segmentation result map. The highlighted weak target region is added to ${M_{th1}}$ in the form of centroid points without affecting the fine segmentation of the strong target region in ${M_{th1}}$. It is worth noting that this strategy exhibits good generality, as it can be seamlessly integrated into existing networks as a post-processing step without requiring any modifications to the original structure.
4 Experiment
4.1 Dataset
Four SIRST datasets (W-IRSTD dataset [39], NUDT-SIRST [29], NUAA-SIRST [25], and IRSTD-1k [33]) are used for experiments. These datasets are introduced as follows:
1) W-IRSTD dataset: The dataset contains 9,000 images, which mainly come from seven public datasets: SIRST-V2 [40], IRSTD-1k [33], IRDST [41], NUDT-SIRST [29], NUDT-SIRST-sea [42], NUDT-MIRSDT [43] and Anti-UAV410 [5]. fig. 3(d) shows some samples from the W-IRSTD dataset. The dataset covers multiple observation perspectives, including land-based, air-based, and space-based. In addition, it includes diverse target types such as extended targets, spot targets, and point targets, spans multiple spectral bands including short-wave infrared, long-wave infrared, and near-infrared, and provides various resolutions such as 256×256, 512×512, and 1024×1024. In the experiment, 7200 samples are used as the training set and 1800 samples are used as the test set.
2) NUAA-SIRST dataset: The real dataset consists of 427 infrared images captured from various scenes, with approximately 10% of the images containing multiple targets. In the experiment, we divide the dataset into training set and test set in a ratio of 213:214. fig. 3(a) shows some samples from the NUAA-SIRST dataset.
3) NUDT-SIRST dataset: The synthetic dataset covers five background types: city, field, highlight, ocean and cloud. About 37% of the images contain two or more targets, and about 32% of the targets are distributed outside the top 10% of the image brightness values. In the experiment, we divide the dataset into training set and test set in a ratio of 663:664. fig. 3(b) shows some samples from the NUDT-SIRST dataset.
4) IRSTD-1k dataset: The real dataset consists of 1001 infrared images with a resolution of 512×512 pixels. It contains small targets of various types, including drones, creatures, ships, and vehicles, with diverse spatial distributions. Meanwhile, the dataset covers a wide range of scenes, such as sky, ocean, and land. In the experiments, we divide the dataset into training set and test set in a ratio of 801:200. fig. 3(c) shows some samples from the IRSTD-1k dataset.
4.2 Experimental settings
1) Experimental environment and parameter settings. The experimental environment is an Ubuntu 18.04 operating system, and the GPU is an RTX 3090 24 GB. The epoch is set to 300 and the learning rate is set to $5e^{-4}$.
Figure 4: ROC curves of models trained with different loss functions.
Table 1: Performance comparison of different losses on the W-IRSTD dataset.
Scheme $IoU$ $P_{d}$ $F_{a}$$10^{-6}$) ↓ Score
SoftIoU Loss [44] 44.08 53.69 12.39 48.89
Dice Loss [45] 50.15 60.09 9.09 55.12
BCE Loss [38] 55.44 59.35 7.02 57.40
EE Loss (ours) 56.26 60.46 8.51 58.36
DM Loss (ours) 56.12 62.45 7.83 59.29
EEDM Loss (ours) 57.30 63.34 14.25 60.32
Table 2: Performance comparison of different scales on the W-IRSTD dataset.
Scheme $IoU$ $P_{d}$ $F_{a}$$10^{-6}$) ↓ Score
256 43.78 48.75 13.79 46.27
384 53.79 64.89 11.19 59.34
512 57.30 63.34 14.25 60.32
640 59.65 72.05 11.55 65.85
768 60.67 70.90 13.53 65.79
896 61.89 71.95 12.32 66.92
1024 60.13 76.17 12.78 68.15
2) Evaluation metrics. This study uses pixel-level evaluation metrics (intersection over union, $IoU$), target-level evaluation metrics (detection rate $P_{d}$ and false alarm rate $F_{a}$) and comprehensive evaluation metrics Score to evaluate the performance of the model. The metric Score is calculated by the $IoU$ and $P_{d}$ under a certain $F_{a}$ constraint. The Score can be used to comprehensively evaluate the region segmentation accuracy and target detection accuracy of the model:
$IoU=\frac{{\sum\nolimits_{i}^{N}{TP[i]}}}{{\sum\nolimits_{i}^{N}{T[i]+p[i]-TP[i]}}}$ (9)
${P_{d}}=\frac{{{T_{correct}}}}{{{T_{All}}}}$ (10)
${F_{a}}=\frac{{{P_{false}}}}{{{P_{all}}}}$ (11)
$Score=\alpha\times IoU+(1-\alpha)\times{P_{d}}$ (12)
where $N$ denotes the number of images in the test set. $TP$ denotes the small target area that the model correctly predicts. $T$ denotes the small target area in the true label. $P$ denotes the small target area predicted by the model. ${T_{correct}}$ denotes the number of correctly predicted targets. ${T_{All}}$ denotes the true target number. ${P_{false}}$ denotes the number of incorrectly predicted target pixels. ${P_{all}}$ denotes the total number of pixels in the image. $\alpha$ is set to 0.5. In addition, $F_{a}$ has a constraint effect. When $F_{a}$ is less than $1e^{-4}$, the final result is considered valid.
4.3 Effect verification of the EEDM loss
To further verify the effect of the proposed EEDM loss, we conduct a detailed experimental exploration, including comparative with other common losses [38, 44, 45] and evaluation of hyperparameter stability. The experiments are conducted on the W-IRSTD dataset. MSDA-Net is selected as the basic model.
1) Comparison with other loss functions: To evaluate the effectiveness of the proposed EEDM Loss, we conduct comparative experiments against commonly used loss functions, including binary cross-entropy Loss (BCE Loss) [38], soft intersection over union Loss (SoftIoU Loss) [44] and Dice Loss [45]. To comprehensively evaluate the effectiveness of the proposed EEDM Loss, we further introduce two ablation variants for comparison: the edge enhancement loss (EE Loss), which contains only the edge enhancement part, and the difficult mining loss (DM Loss), which contains only the difficult pixel mining part. This design enables a quantitative analysis of the respective contributions of edge enhancement and difficult sample mining to the overall detection performance.
From table 2, compared with other losses, the proposed EEDM loss results in more obvious improvements in both $IoU$ and $P_{d}$, verifying its superiority in infrared small target detection tasks. In contrast, the widely used SoftIoU Loss exhibits relatively inferior performance on the complex and diverse W-IRSTD dataset. At the same time, compared with BCE Loss, EE Loss achieves improvements of 1.48% in $IoU$ (from 55.44 to 56.26), 1.87% in $P_{d}$ (from 59.35 to 60.46), and 1.67% in Score (from 57.40 to 58.36), indicating that incorporating edge features during training can effectively enhance detection performance. In addition, compared with BCE Loss, DM Loss achieves improvements of 1.23% in $IoU$ (from 55.44 to 56.12), 5.2% in $P_{d}$ (from 59.35 to 62.45), and 3.19% in Score (from 57.40 to 59.29), indicating the effectiveness of difficult pixel mining in this task. From fig. 4, the ROC curve of the model trained with EEDM Loss consistently stays above the others, achieving the highest AUC, which indicates superior performance across all thresholds compared to other loss functions.
Figure 5: Performance analysis of the hyperparameters $w$ and $p$ on the W-IRSTD dataset. From left to right, the figures show the model performance in terms of $IoU$, $P_{d}$, $F_{a}$, and Score under different combinations of $(w,p)$. Red indicates the optimal value in each metric.
Table 3: Comparison of the fusion performance of multi-scale models on the W-IRSTD dataset.
Scheme $IoU$ $P_{d}$ $F_{a}$$10^{-6}$) ↓ Score
896 61.89 71.95 12.32 66.92
896+768 63.12 70.85 9.32 66.99
896+768+1024 64.49 72.58 9.17 68.54
896+768+1024+640 64.11 72.42 7.96 68.27
896+768+1024+640+512 63.60 70.43 7.25 67.02
896+768+1024+640+512+384 62.83 68.72 6.75 65.78
896+768+1024+640+512+384+256 61.44 65.94 6.17 63.69
Table 4: Performance comparison of adjusting the threshold th1 on the W-IRSTD dataset.
Scheme th1 $IoU$ $P_{d}$ $F_{a}$$10^{-6}$) ↓ Score
896+768+1024 (0.5) 0.5 64.49 72.58 9.17 68.54
896+768+1024 (0.45) 0.45 65.53 74.86 10.51 70.20
896+768+1024 (0.4) 0.4 65.91 76.44 13.03 71.18
896+768+1024 (0.35) 0.35 65.58 78.56 15.67 72.07
896+768+1024 (0.3) 0.3 64.22 80.29 20.57 72.26
896+768+1024 (0.25) 0.25 61.83 82.34 26.85 72.09
896+768+1024 (0.2) 0.2 58.34 83.91 35.93 71.13
896+768+1024 (0.15) 0.15 53.43 85.28 50.52 69.36
896+768+1024 (0.1) 0.1 46.27 86.43 76.12 66.35
896+768+1024 (0.05) 0.05 35.15 87.72 138.46
Table 5: Performance comparison of adjusting the threshold th2 on the W-IRSTD dataset.
Scheme th1 th2 $IoU$ $P_{d}$ $F_{a}$$10^{-6}$) Score
896+768+1024 (0.3) 0.3 - 64.22 80.29 20.57 72.26
896+768+1024 (0.3+0.5) 0.5 63.96 81.61 21.26 72.79
896+768+1024 (0.3+0.4) 0.4 63.78 83.02 21.76 73.40
896+768+1024 (0.3+0.3) 0.3 63.45 85.28 22.66 74.37
896+768+1024 (0.3+0.2) 0.2 62.81 87.51 24.29 75.16
896+768+1024 (0.3+0.1) 0.1 61.42 89.98 28.11 75.70
896+768+1024 (0.3+0.05) 0.05 59.44 91.05 33.05 75.25
2) Validation of hyperparameter stability: To comprehensively evaluate the performance of the proposed EEDM Loss, we conduct detail experiments on two key hyperparameters (the edge weight coefficient $w$ and the difficult pixel mining ratio $p$) to investigate the model’s sensitivity and robustness to different hyperparameter settings. Specifically, the edge weight coefficient $w$ is set to 1, 3, 4, 5, 7, and the difficult pixel mining ratio $p$ is set to 0.1, 0.3, 0.5, 0.7, 0.9.
From fig. 5, the proposed EEDM Loss can maintain relatively stable performance across different hyperparameter configurations. This indicates that the loss function is robust to the hyperparameters $w$ and $p$, as it consistently enhances model performance within a reasonable range of these hyperparameters. Based on the experimental results, the metrics Score reaches the optimal value when $w$ is 4 and $p$ is 0.5. Specific analysis shows that when the $w$ is too small, the model pays insufficient attention to edge information, leading to decreases in $IoU$ and $P_{d}$. When $w$ is too large, the model will overly rely on edge features, thereby neglecting semantic and other critical features, which negatively impacts the overall detection performance. It is worth noting that when $w$ is extremely reduced to 1, the EEDM Loss degenerates into the DM Loss. Meanwhile, when the $p$ is too small, the network tends to excessively focus on difficult pixels during training, which may hinder its generalization ability and lead to degraded performance. Conversely, when $p$ is too large, the mining of difficult pixels will be insufficient, thereby weakening the model’s adaptability in complex scenarios. It is worth noting that when $p$ increases to 1, the EEDM Loss degenerates into the EE Loss.
4.4 Effect verification of the multi-scale fusion
To evaluate the effectiveness of the multi-scale fusion strategy, we conduct extensive experiments on the W-IRSTD dataset, using MSDA-Net as the basic model and EEDM Loss as the loss function. Specifically, we resize the original images to various resolutions, including 256×256, 384×384, 512×512, 640×640, 768×768, 896×896, and 1024×1024, to systematically explore the impact of input scale on model performance. Based on this, we take the best-performing single-scale model as the foundation and progressively integrate suboptimal models to further evaluate the effectiveness of the multi-scale fusion strategy. For clarity, each model is named based on its corresponding input resolution. For example, the model trained with 256×256 images is denoted as the 256 model.
From table 2, the 1024 model achieves the best performance. Compared with the worst performing 256 model, the 1024 model improves the $IoU$, $P_{d}$, and Score by 37.35% (from 43.78 to 60.13), 56.25% (from 48.75 to 76.17) and 47.29% (from 46.27 to 68.15), respectively. The analysis reveals that when images from the W-IRSTD dataset are downsampled to smaller resolutions, such as 256×256, the model performance drops significantly. This degradation is primarily attributed to the loss of pixel-level information during the downsampling process, which causes small targets to shrink or even vanish entirely, making it difficult for the network to extract effective features. In addition, considering that the proposed AS strategy will be used later, it can effectively improve the detection rate of the model. Therefore, to ensure the segmentation accuracy, we choose the 896 model with the best $IoU$ metrics as the basis model, and fuse the suboptimal scale models in the $IoU$ metrics one by one. The experimental results are shown in table 5.
From table 5, the fusion of the 896, 768, and 1024 models yields the most significant performance improvement. Specifically, compared with the 896 model, which has the best performance at a single scale, the fusion of the 896 model, the 768 model and the 1024 model can improve its $IoU$, $P_{d}$ and Score by 4.20% (from 61.89 to 64.49), 0.88% (from 71.95 to 72.58) and 2.42% (from 66.92 to 68.54), respectively. At the same time, we find that when the performance of the fused scale models degrades severely, it will lead to the degradation of the performance of the final model. In addition, the multi-scale fusion model will perform better on $F_{a}$, which shows that the multi-scale fusion strategy is conducive to reducing the missed detections and further improving the reliability of the model. To more intuitively demonstrate the effect of multi-scale model fusion, we show the prediction results of the 896 model, 768 model, 1024 model and their fusion model in fig. 6. It is worth noting that although multi-scale model fusion significant improvements in detection performance, it inevitably increases inference time. Therefore, a reasonable trade-off between accuracy and efficiency should be made based on specific requirements in practical applications to achieve optimal overall performance.
Figure 6: Detection results of the fusion model and its corresponding single model. From top to bottom, each row sequentially displays the original image, detection results from the 768 model, 896 model, 1024 model, and the fused model, followed by the corresponding ground truth. In the visualizations, red denotes the correct detections, green denotes the false detections, and yellow denotes missed detections.
Figure 7: 2D comparison of the effects of the AS strategy. In each group of images, the ones with green borders represent detection results without the AS strategy, while the ones with red borders represent detection results after applying the AS strategy. Local zoom has been applied to emphasize the differences in key areas.
4.5 Effect verification of the AS strategy
To verify the AS strategy, we adjust the thresholds th1 and th2 on the optimal multi-model fusion scheme (“896+768+1024”). Table 5 shows the experimental results obtained by adjusting the threshold th1 alone without adjusting the threshold th2. Table 5 shows the experimental results obtained by adjusting the threshold th2 on the basis that th1 is set to the optimal value in table 5.
From table 5, when the threshold th1 is set to 0.3, the model achieves the best overall performance. Compared to the conventional setting of th1 at 0.5, adjusting th1 to 0.3 leads to 5.43% (from 68.54 to 72.26) improvement in the comprehensive metric Score. The key to this improvement lies in the model’s ability to significantly enhance the detection rate $P_{d}$ without notably affecting segmentation accuracy. Specifically, $P_{d}$ increases by 10.62% (from 72.58 to 80.29). Notably, when the threshold th1 is reduced to 0.05, $F_{a}$ will surge and exceed the constraint range.
From table 5, the introduction of th2 significantly improves the detection rate of the network while maintaining segmentation accuracy for infrared small targets and effectively controlling the false alarm rate. On the W-IRSTD dataset, based on the multi-scale fusion scheme of “896+768+1024”, the model achieves optimal performance when th1 is set to 0.3 and th2 is set to 0.1. Specifically, compared to using only th1 set to 0.3, the further introduction of th2 set to 0.1 leads to a 4.76% (from 72.26 to 75.70) improvement in Score metrics.
To more intuitively demonstrate the effectiveness of the AS strategy, we present some results as shown in fig. 7. For extended targets, the appropriate application of the AS strategy helps refine the shape and details of the target. For point or spot targets, the strategy helps improve detection rates and effectively capture weak target features.
Table 6: $IoU$ (%), $P_{d}$ (%) and $F_{a}$$10^{-6}$) values of different methods achieved on the separate NUAA-SIRST, NUDT-SIRST, and IRSTD-1k datasets. (213:214), (663:664) and (800:201) denote the division of training samples and test samples. Red denotes that the $F_{a}$ exceeds the acceptable upper limit, making the results for this group invalid. Green denotes the optimal result.
NUAA-SIRST (213:214) NUDT-SIRST (663:664) IRSTD-1k (800:201)
Scheme Description $IoU$ $P_{d}$ $F_{a}$$10^{-6}$) ↓ Score $IoU$ $P_{d}$ $F_{a}$$10^{-6}$) ↓ Score $IoU$ $P_{d}$ $F_{a}$$10^{-6}$) ↓ Score
Original 62.24 94.30 111.84 - 65.36 96.08 37.48 80.72 61.08 93.60 22.51 77.34
+LW-FEST 71.28 96.96 42.94 84.12 67.10 97.46 17.53 82.28 63.43 94.61 38.41 79.02
ACM [25] +FEST 73.57 97.72 34.78 85.65 73.16 98.52 9.86 85.84 64.77 95.96 33.54 80.37
Original 66.96 93.54 67.83 80.25 70.84 97.35 24.40 84.10 62.93 92.59 29.87 77.76
+LW-FEST 70.19 97.34 38.42 83.77 71.03 97.88 20.63 84.46 64.26 95.62 15.68 79.94
ALCNet [26] +FEST 74.28 98.10 28.95 86.19 74.91 98.94 17.35 86.93 66.85 97.31 13.27 82.08
Original 72.87 94.30 37.79 83.59 90.65 98.84 9.93 94.75 65.23 93.94 18.11 79.59
+LW-FEST 73.24 96.20 50.83 84.72 92.54 99.15 9.79 95.85 65.70 94.95 16.25 80.33
MLCL-Net [27] +FEST 73.61 96.20 34.78 84.91 92.70 99.37 10.32 96.04 66.20 95.62 18.09 80.91
Original 75.29 95.44 36.84 85.37 91.52 98.41 5.84 94.97 66.91 93.94 29.68 80.43
+LW-FEST 75.71 96.20 23.94 85.96 92.65 99.47 10.34 96.06 67.47 95.96 34.92 81.72
ALCL-Net [28] +FEST 77.03 96.96 18.87 87.00 92.86 99.58 6.18 96.22 67.47 97.64 22.72 82.56
Original 76.25 96.20 29.56 86.23 91.14 98.73 10.50 94.94 68.69 94.61 43.49 81.65
+LW-FEST 77.40 97.72 26.34 87.56 92.25 99.37 18.13 95.81 69.43 95.62 38.32 82.53
DNA-Net [29] +FEST 79.06 98.10 24.97 88.58 94.20 99.47 6.99 96.84 70.09 96.30 26.97 83.20
Original 78.77 98.10 31.53 88.44 92.91 99.15 2.57 96.03 68.40 93.60 17.23 81.00
+LW-FEST 78.82 98.48 22.43 88.65 93.34 99.47 3.13 96.41 67.77 94.95 23.57 81.36
GGL-Net [31] +FEST 79.59 98.10 17.70 88.85 93.39 99.58 3.79 96.49 70.02 95.29 25.89 82.66
Original 77.08 96.20 46.50 86.64 92.04 98.31 14.02 95.18 70.70 94.26 31.79 82.48
+LW-FEST 77.29 96.20 31.69 86.75 93.21 99.58 9.15 96.40 72.36 95.62 33.04 83.99
UIUNet [30] +FEST 79.37 96.20 29.98 87.79 93.71 99.58 4.04 96.65 72.86 96.97 29.57 84.92
Original 77.43 96.20 42.10 86.82 93.21 98.73 2.94 95.97 71.03 93.94 20.10 82.49
+LW-FEST 78.27 98.10 20.17 88.19 94.29 99.37 1.82 96.83 70.89 96.00 28.87 83.45
MSDA-Net [32] +FEST 80.51 97.72 15.23 89.12 94.39 99.68 3.13 97.04 71.19 97.98 25.07 84.59
Figure 8: Visualization of several SIRST methods on the NUAA-SIRST dataset. Red denotes the correct detections, blue denotes the false detections, and yellow denotes the missed detections. Every two rows from top to bottom: Image, Original SIRST Network, SIRST Network equipped with FEST framework, True label.
4.6 Validation of the FEST Framework on multiple datasets
To validate the effectiveness and robustness of the proposed FEST framework, we integrate multiple advanced SIRST networks into the framework and conduct experiments on three public datasets. To ensure a fair comparison, all models are retrained under the same settings, with the number of training epochs uniformly set to 300 to reduce training costs. This setup differs from that used in our previous work. In addition, we explore the performance gain brought by using only EEDM Loss and AS strategy in the FEST framework, with almost no additional resource consumption. This lightweight variant is referred to as the LW-FEST framework. The thresholds th1 and th2 in both the FEST and LW-FEST frameworks are set to 0.5 and 0.1, respectively. Notably, considering the characteristics of each dataset, we adopt differentiated configurations in the multi-scale fusion strategy: for NUAA-SIRST and IRSTD-1k, we fuse models with sizes of 512 and 768, whereas for NUDT-SIRST, we fuse models with sizes of 256 and 384.
Figure 9: Visualization of several SIRST methods on the NUDT-SIRST dataset. Red denotes the correct detections, blue denotes the false detections, and yellow denotes the missed detections. Every two rows from top to bottom: Image, Original SIRST Network, SIRST Network equipped with FEST framework, True label.
Figure 10: Visualization of several SIRST methods on the IRSTD-1k dataset. Red denotes the correct detections, blue denotes the false detections, and yellow denotes the missed detections. Every two rows from top to bottom: Image, Original SIRST Network, SIRST Network equipped with FEST framework, True label.
From table 6, embedding various SIRST networks into the lightweight LW-FEST framework can effectively improve the model performance without increasing resource consumption. These results demonstrate that the LW-FEST framework exhibits strong robustness and generalization capabilities. It can be seamlessly integrated with various SIRST networks, effectively enhancing their feature extraction and representation abilities. Meanwhile, when networks with relatively poor performance (such as ACM) are embedded into the framework, the performance improvement is more significant. In addition, the application of the complete FEST framework can further improve the detection performance. Specifically, on the NUAA-SIRST dataset, compared with the original SIRST network, the SIRST network equipped with the FEST framework achieves a performance improvement of 0.46% - 6.89% in the evaluation metric Score. On the NUDT-SIRST dataset, compared with the original SIRST network, the SIRST network equipped with the FEST framework achieves a performance improvement of 0.48% - 6.34% in the evaluation metric Score. On the IRSTD-1k dataset, compared with the original SIRST network, the SIRST network equipped with the FEST framework achieves a performance improvement of 1.66% - 5.56% in the evaluation metric Score.
To further demonstrate the superiority of the proposed framework, we present visual comparisons of partial difficult samples detection results across three datasets, as shown in fig. 8, fig. 10, and fig. 10. Equipping SIRST networks with the FEST framework leads to clearer target edges, more comprehensive detection results, and enhanced perception and localization accuracy for small targets in diverse detection scenarios.
5 Conclusion
This manuscript proposes a feature-enhanced and sensitivity-tunable (FEST) framework, aiming to provide an efficient and universal optimization framework for infrared small target detection tasks. Specifically, we propose an edge enhancement difficulty mining (EEDM) loss, which guides the network to focus more on challenging regions and edge features during training, thereby promoting fine-grained learning and positive optimization of the model. Meanwhile, we adopt a multi-scale fusion strategy, which enhances the network’s perception and adaptability to multi-scale features of multi-size targets by integrating feature information from different scales. In addition, to fully utilize the probability mask information, we design an adjustable sensitivity (AS) strategy for network post-processing. This strategy employs a tunable dual threshold mechanism, which significantly improves the detection rate of infrared small targets while maintaining the segmentation accuracy of the model. Extensive experimental results demonstrate that the proposed FEST framework achieves significant performance gains and exhibits strong robustness on multiple public datasets. Notably, the lightweight LW-FEST framework achieves notable performance gains with minimal additional resource consumption.